OpenStudy (carolina4567):

Eva played her favorite video game for 14 hours last week. Today, Eva's parents restricted her to 4 hours each week for the next 7 weeks. Create the function, f(x), that models Eva's total video game time, and explain what each number in the situation represents graphically, using complete sentences.

OpenStudy (carolina4567):

@Loser66

OpenStudy (carolina4567):

Help Me

OpenStudy (carolina4567):

I need to create an F(x) function :(

OpenStudy (dangerousjesse):

You're looking for the total, so it'll be \(F(x)=new~ time ~allowed\times ~number ~of ~weeks+ old ~time~ allowed\times~ previous~ week\)

OpenStudy (dangerousjesse):

So \(\textsf{F(x)=4x+14}\)

OpenStudy (carolina4567):

Oh ok Thank u so much @DangerousJesse umm i have two more questios I hope u dont mind

OpenStudy (dangerousjesse):

That's fine :)

OpenStudy (carolina4567):

Pedro has created the function f(x) = the quantity of 4x minus 3, divided by 2 to represent the number of assignments he has completed, where x represents the number of weeks in the course. Pedro discovers that, using the inverse function to solve for x = 30, he can predict when he will have 30 assignments completed. Explain to Pedro how to accomplish this, using complete sentences.

OpenStudy (carolina4567):

@DangerousJesse

OpenStudy (carolina4567):

@phi

OpenStudy (phi):

it's hard to read your equations when written out (which is why people invented the way we do it) is it \[ f(x) = \frac{4x-3}{2} \] ?

OpenStudy (carolina4567):

Yes

OpenStudy (phi):

the technique to find the inverse function is write it as \[ y = \frac{4x-3}{2} \] then "swap x and y" to get \[ x = \frac{4y-3}{2} \] now "solve for y" The first step is multiply both sides by 2 can you do that ?

OpenStudy (carolina4567):

2 x X = 4y - 3 / 2 x 2

OpenStudy (carolina4567):

is it like tha

OpenStudy (phi):

almost. we leave out the multiply sign x (because it will be confused with the variable x) so it would look like this \[ 2x = \frac{4y-3}{2} \cdot 2 \] when you have 2/2 (on the right side), that simplifies to 1, so you have \[ 2x = 4y-3 \] now add +3 to both sides. what do you get ?

OpenStudy (carolina4567):

2x = 4y - 3 would it be like this +3 +3

OpenStudy (phi):

ok, but write it on one line 2x+3 = 4y-3 + 3 can you simplify the right side?

OpenStudy (carolina4567):

Ok would it be 5x = 4y - 6 Im kinda confsed im sorry

OpenStudy (phi):

2x + 3 stays like that. You have 2 x's plus a 3 on the right side you have 4 y's , a 3 and take away a 3

OpenStudy (carolina4567):

Oh ok so then 2x + 3 = 4y

OpenStudy (carolina4567):

is that correct @phi

OpenStudy (phi):

yes, much better now divide both sides by 4: \[ \frac{2x + 3 }{4}= \frac{4y }{4} \]

OpenStudy (phi):

can you simplify the right side?

OpenStudy (carolina4567):

Ok so would it be 4 times 2x + 3 / 4 = 4y/4 times 4

OpenStudy (phi):

we don't want to multiply both sides by 4. Here is the idea: we want "y" by itself when we have \[ \frac{4y}{4}\] that is the same as \[ \frac{4}{4} \cdot \frac{y}{1} \] If you multiply fractions, you multiply top time top and bottom times bottom we would get back 4y/4 However, 4/4 is 1, so \[ \frac{4}{4} \cdot \frac{y}{1} \\ 1 \cdot \frac{y}{1} \\ 1\cdot y\\y \]

OpenStudy (phi):

in other words, the reason we divided 4y by 4 is to make it y (by itself) to keep things even, we divide the other side by 4 we get \[ \frac{2x+3}{4} = y\] now you use this inverse function to answer the question. replace x with 30 can you do that ?

OpenStudy (carolina4567):

ok

OpenStudy (carolina4567):

2(30) + 3 / 4 = y 60 + 3 / 4 = y 63 / 4 = y 15.75

OpenStudy (phi):

yes. 15.75 is the number of weeks it will take to finish 30 assignments.

OpenStudy (carolina4567):

EEEKK THANK U SO MUCH :)

OpenStudy (phi):

I hope you have time to review how to "solve" equations. The steps (once you learn them) are easy to do.